OpenStudy (anonymous):

Solve the system of equations using the elimination method. Enter the answer as an ordered pair. -x + y = -3 3x + 2y = 9

hero (hero):

Hint: 1. Multiply the first equation by 3 2. Add both equations together. 3. Solve for y

OpenStudy (anonymous):

@Hero How Would I Set Up The Equations ? I Suck At This :((

hero (hero):

Alternatively, you can put both equations in the form y = mx + b, then set y = y

OpenStudy (anonymous):

@Hero so the first equation will be -x+y=-3*3?

OpenStudy (cunningham2000):

3(-x+y=3) would be -3x + 3y = 9 Multiply each term by three

OpenStudy (cunningham2000):

Sorry -3x + 3y = -9 I dropped the negative

OpenStudy (cunningham2000):

When you put the first equation over the second equation your x term will 'eliminate' or become zero. then you can solve for y.

OpenStudy (anonymous):

@cunningham2000 -3x+3y=-9 3x+2y=9 so it equals to x=0 right ?